how to find heat loss

Grade:11

1 Answers

Aman Bansal
11 years ago

Dear Paras,

alculating a heat load is necessary before the installation of a radiant heating system can begin, since different types of radiant heating systems have different BTU output values.


A typical heat load calculation consists of surface heat loss calculation and heat loss due to air infiltration. Both should be done separately for every room in the house, so having a floor plan with dimensions of all walls, floors, ceiling, as well as doors and windows is a good place to start.



Below is a sample 6-step manual to surface heat loss calculation:


Step 1 – Calculate Delta T (Design Temperature):


Delta T is a difference between indoor design temperature (T1) and outdoor design temperature (T2), where indoor design temperature is typically 68-72F depending on your preference, and outdoor design temperature is a typical low during the heating season. The former can be obtained by calling your local utility company.


Assuming that T1 is 72F and T2 is –5F,  Delta T = 72F - (-5F) = 72F + 5F = 77F



Step 2 – Calculate surface area:


If the calculation is done for an outside wall, with windows and doors, the calculations for the window and door heat loss should be done separately.


Wall Area = Height x Width - Door Surface - Window Surface


Wall Area = 8ft x 22ft - 24sq ft - 14sq ft = 176sq ft - 38sq ft = 138sq ft



Step 3 – Calculate U-value:


Use “Typical R-values and U-values” guide to obtain the wall R-value.


U-value = 1 / R-value U-value = 1 / 14.3 = 0.07



Step 4 – Calculate wall surface heat loss:


Surface heat loss can be calculated using the formula below:


Surface Heat Loss = U-value x Wall Area x Delta T


Surface Heat Loss = 0.07 x 138sq ft x 77F = 744 BTUH

(U-value is based on assuming a 2x4 wood frame wall with 3.5” fiberglass insulation)



Step 5 – Calculate total wall heat loss:


Follow the steps 1 through 4 to calculate heat loss separately for windows, doors, and ceiling.


Door Heat Loss = 0.49 x 24sq ft x 77F = 906 BTUH

(U-value is based on assuming a solid wood door)


Window Heat Loss = 0.65 x 14sq ft x 77F = 701 BTUH

(U-value is based on assuming a double-panel window)


Ceiling Heat Loss = 0.05 x 352sq ft x 77F = 1355 BTUH

(U-value is based on assuming a 6” fiberglass insulation. Ceiling surface is 22ft x 16ft)


Now, add all the number together.

 

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert